# finantial support acknowledgement

### v2

@@ VI-A2 Targeted Backdoor Attacks
Notably, Figure 4 illustrates how FedCC significantly nullifies backdoor confidence. Compared to oscillating values of other methods, the stable low confidence values further emphasize FedCC’s resilience and independence from the specific data distribution across clients. This advantage stems from its effective utilization of CKA, enabling the calculation of similarity values even when models are trained on different datasets with varying data distributions. Whereas it is true that other baselines, such as Bulyan, also reduce confidence, we argue that FedCC not only preserves but potentially increases accuracy by precisely discriminating against malicious clients.
VI-B IID Data Environment
@@
VI-B1 Untargeted Model Poisoning Attacks
The bottom section of Table III (fifth to seventh rows) provides valuable insights into the test accuracy of a global model trained on three datasets under untargeted attacks in IID settings. Although the trends observed in non-IID experiments appear similar, the accuracy values are slightly higher in this case due to the identical nature of the data.
Under the Fang-Krum attack, we can observe relatively low accuracy values for Krum and Bulyan, as the attack explicitly targets Krum and its variations. However, when Coomed is applied, the accuracy increases from 75.55% (FedAvg) to 79.16% (Coomed), indicating that coordinate-wise median values closely align with the weights of the global model with minimal deviation. FLARE’s low accuracy is due to the continuous scaling of weights even when the model is already optimized. In contrast, FedCC does not scale weights but focuses solely on optimizing the model to the minimum by averaging chosen clients. As a result, FedCC achieves the highest accuracy (89.23%, 67.12%, and 14.51% for each dataset) by accurately identifying and filtering out compromised clients.
Under the Fang-Med attack, the accuracy of FedAvg experiences a significant drop (89.89% to 20.86%) due to the peculiarity of outliers when the clients have IID data. Notably, Krum, Coomed, and Bulyan demonstrate similar accuracy values, indicating that coordinate-wise median values indeed represent benign clients’ parameters. Meanwhile, FedCC consistently achieves the highest accuracy, with values of 88.44%, 64.74%, and 17.83% for each dataset. Notably, the remarkable growth observed in experiments with CIFAR100, with a 66% and 58.7% increment under Fang-Krum and Fang-Med attacks, respectively, surpasses the second-highest improvement. These results strongly validate the effectiveness of FedCC in mitigating untargeted model poisoning attacks in a simple IID setting.
VI-B2 Targeted Backdoor Attacks
The last row of Table III and the bottom three figures of Figure 4 are the test accuracy and backdoor confidence under targeted backdoor attacks in an IID setting. In targeted backdoor attacks, the global model is covertly trained to misclassify a specific attacker-chosen class to an attacker-chosen label while leaving the remaining classes intact. As a result, the test accuracy has remained relatively high across the different methods.
However, a critical observation is the remarkable confidence reduction achieved by FedCC. The backdoor confidence with FedCC approaches zero, effectively neutralizing the backdoor task. In comparison, the attack confidence of FedAvg and Bulyan, despite exhibiting higher test accuracy for CIFAR10, remains higher than that of FedCC. Note that the confidence of the other methods fluctuates by a large margin. Specifically, FLARE’s backdoor confidence oscillates even though it occasionally records lower confidence than ours (e.g., CIFAR10). This means that its performance is less stable than that of FedCC, which consistently minimizes backdoor confidence. It is crucial to emphasize that reducing attack confidence is as significant as maintaining the main test accuracy since the attacker’s objective is to train the model on backdoor tasks stealthily, but the defender’s goal is to nullify them. Hence, our proposed method achieves the highest accuracy for fMNIST and CIFAR100 and comparable accuracy for CIFAR10 while effectively nullifying the backdoor confidence. This outcome underscores the effectiveness and robustness of our approach in mitigating targeted backdoor attacks.
VII Performance in Varying FL Settings
@@
We additionally measure the effectiveness of FedCC in various non-IID settings, including different numbers of malicious clients, fractions of participation, and numbers of local epochs. Since Coomed was as effective as ours, and FLARE also used PLRs, we compare the three methods to FedCC, including FedAvg.
VII-A Impact of the Number of Malicious Clients
@@
Figure 5: Accuracy with different numbers of malicious clients
@@
We assess the impact of different numbers of malicious clients under untargeted attacks, specifically Fang-Krum and Fang-Med. In this set of experiments, we fix the total number of clients and participation fraction at 10 and 1, respectively. To adhere to the assumption that the number of attackers is less than half of the participating clients, we investigate the test accuracy considering scenarios involving up to four malicious clients.
@@
Figure 5 illustrates the accuracy in the given experimental settings, and it is evident that FedCC outperforms the other defense methods. We observe a general trend where the accuracy of defense methods decreases as the number of attackers increases. Notably, Coomed experiences the most significant drop in accuracy when transitioning from one to four malicious clients. This can be attributed to the fact that as the participation of malicious clients increases, there is a higher likelihood that the median values will be influenced by the malicious contributions.
@@
VII-B Impact of Fraction
TABLE IV: Accuracy under untargeted attacks with different fractions of clients being selected. fM refers to fMNIST, C10 refers to CIFAR10, Med refers to Coordinate wise median.
Fang-Krum Fang-Med
@@ VII-B Impact of Fraction
We examine the impact of the participating fraction of clients where the number of total clients is fixed at 100. The fractions are 0.1, 0.3, and 0.5, such that the numbers of clients being aggregated are 10, 30, and 50, respectively. The proportion of malicious clients is fixed at 0.3, such that the malicious clients are 3, 9, and 15, respectively. Table IV summarizes the accuracy. We can observe that the accuracy tends to increase as the participation fraction is greater under Fang-Krum attacks. It implies that the more client participation in a non-IID setting, the more accurate the global model is if the attacks are mitigated. Under Fang-Med attacks, the accuracy is the highest with FedCC and second highest with Coomed. Throughout the experiment, FLARE prevented the global model from convergence, and the test accuracy fluctuated. The fluctuation is presumably due to the weight scaling while randomly selecting clients. Distinct from previous experiments, the server chooses a fraction of the client every round, meaning not-yet-trained models could be selected. In this environment, weight scaling leveraged by FLARE constantly improperly scales the local weights, causing divergence and fluctuation. These experimental results indicate that selectively choosing clients based on CKA similarity is more reliable than taking the median value of each coordinate in a non-IID setting. We also observed that as the participation fraction grows, the accuracy grows. This is due to the rising number of aggregated local models; the more models are aggregated, the more general global model is generated.
VII-C Impact of the Number of Local Epochs
TABLE V: Accuracy under Backdoor Attack with doubled numbers of local epochs
Targeted (Diff Local Epoch)
Dataset FedAvg Coomed FLARE FedCC
fMNIST 69.40 72.11 72.32 82.31
CIFAR10 53.85 49.18 21.61 54.94
Figure 6: Backdoor confidence when the numbers of local epochs are doubled.
To examine the impact of the number of local epochs, we doubled up the number of local epochs such that the local models are trained for six epochs in benign clients and twelve epochs in malicious clients. The main task accuracy and confidence are summarized and illustrated in Table V and Figure 6. As expected, increasing the number of local epochs would not affect the accuracy or confidence but bring the convergence forward while being better at maintaining test accuracy. It is true that Coomed and FLARE significantly reduce the backdoor confidence, as well as FedCC; however, the test accuracy has even increased with FedCC due to its precise discrimination against malicious clients.
VIII Conclusion and Future Work
FL has emerged in response to the rising concerns about privacy breaches while employing AI techniques. FL, however, is vulnerable to model poisoning attacks due to a server’s blindness to local datasets, making it challenging to assume data distribution and model parameter integrity. In line with this, existing robust aggregation algorithms and defense approaches are orthogonal to various attacks and lack consideration of non-IIDness. Throughout exhaustive experiments, FedCC mitigates both untargeted and targeted (or backdoor) attacks while demonstrating its effectiveness in non-IID data environments. We leave theoretical guarantees of FedCC and experiments when clients have distinct neural network architectures or tasks as future work.
@@ VIII Conclusion and Future Work
VIII-A Dendrograms to Measure Cluster Distance
Figure 7: Dendrogram in a non-IID setting
Figure 7 illustrates the distance between two clusters of clients using dendrograms with a single linkage metric and correlation distance method. The first two and last columns correspond to untargeted (Fang-Krum, Fang-Med) and targeted attacks, respectively. The total number of clients is set to 10; the participation fraction is 1.0. The number of malicious clients is 3 for untargeted and 1 for a targeted attack, with fedAvg. Observably, the height difference between the root of clusters is maximum in the second last layer (PLR).
VIII-B Model Architectures
The tables below summarize model architectures for each benchmark dataset (fMNIST, CIFAR10, CIFAR100). Considering that the end devices are generally incapable of handling heavy computation due to resource or communication constraints, we used lightweight models for the experiments.
TABLE VI: CNN Architecture for fMNIST dataset
Layer In Out Ker / Str / Pad Activation
conv2d_1 1 64 $5\times 5$ / 1 / 0 ReLU
conv2d_2 64 64 $5\times 5$ / 1 / 0 ReLU
dropout - - 0.25 -
flatten - 25600 - -
fc_1 25600 128 - -
dropout - - 0.5 -
fc_2 128 10 - -
TABLE VII: CNN Architecture for CIFAR10 dataset
Layer In Out Ker / Str / Pad Activation
conv2d_1 3 64 $3\times 3$ / 1 / 0 ReLU
maxpool2d $2\times 2$ / - / 0 -
conv2d_2 64 64 $3\times 3$ / 1 / 0 ReLU
maxpool2d $2\times 2$ / - / 0 -
conv2d_3 64 64 $3\times 3$ / 1 / 0 ReLU
maxpool2d $2\times 2$ / - / 0 -
flatten - 256 - -
dropout - - 0.5 -
fc_1 256 128 - -
fc_2 128 10 - -
TABLE VIII: LeNet Architecture for CIFAR100 dataset
Layer In Out Ker / Str / Pad Activation
conv2d_1 3 6 $5\times 5$ / 1 / 0 ReLU
maxpool2d $2\times 2$ / 2 / 0 -
conv2d_2 6 16 $5\times 5$ / 1 / 2 ReLU
maxpool2d $2\times 2$ / 2 / 0 -
flatten - $4$
fc_1 44944 120 - ReLU
fc_2 120 84 - ReLU
fc_3 84 100 - -
Acknowledgment

### v3

@@ 6.1.2 Targeted Backdoor Attacks
Notably, Figure 3 demonstrates that FedCC significantly reduces backdoor confidence. Note that since the backdoor confidence of DBA fluctuates a lot due to its distributed nature, we omit it for brevity. Unlike other methods, such as FLTrust or multi-Krum, which exhibit fluctuating confidence values, FedCC maintains consistently low confidence, emphasizing its resilience and independence from client-specific data distributions. This advantage stems from the effective utilization of CKA, enabling similarity calculations even when models are trained on different datasets with varying distributions. While other baselines, like Bulyan or FLTrust, also reduce confidence, we argue that FedCC not only preserves but potentially increases accuracy by precisely distinguishing malicious clients.
6.2 IID Data Environment
@@
Table 4: Test Accuracy under untargeted attacks in IID setting.
Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
Fang -Krum IID fM 75.55 31.66 87.78 87.62 50.30 89.53 79.16 89.57
C10 49.67 40.86 63.42 57.40 12.67 68.25 25.77 69.84
C100 13.72 1.04 7.64 6.17 1.59 17.14 7.49 18.47
Fang -Med IID fM 20.86 85.33 89.53 86.70 87.45 21.36 71.08 89.66
C10 9.51 54.28 69.68 59.20 57.69 9.92 49.82 70.52
C100 0.87 12.27 16.52 14.43 12.56 1.16 5.83 17.83
6.2.1 Untargeted Model Poisoning Attacks
Table 4 shows the test accuracy of the global model trained on three datasets under untargeted attacks in IID settings. While trends in these experiments are similar to those in the non-IID setting, the accuracy values are slightly higher due to the more consistent data distribution across and updates from clients in the IID case. A notable deviation from the non-IID experiments is when Coomed is applied: the accuracy increases from 75.55% (FedAvg) to 87.62% (Coomed). This indicates that the coordinate-wise median values closely align with the global model’s weights with minimal deviation and represent benign models well, as the adversary’s impact is limited due to the majority of benign clients. FedCC achieves the highest accuracy across all methods (89.57%, 69.84%, and 18.47% for each dataset) due to its effective CKA similarity measurement and sophisticated layer-wise aggregation strategy.
Under the Fang-Med attack, the accuracy of FedAvg experiences a significant drop (89.89% to 20.86%) due to the peculiarity of outliers when the clients have IID data. Similar to the result of Fang-Krum, Krum, Coomed, and Bulyan demonstrate similar accuracy values, indicating that coordinate-wise median values indeed represent benign clients’ parameters. FLTrust, however, is less effective in mitigating the Fang-Med attack than Fang-Krum, even in the IID setting; it then means that the performance difference is not due to data distribution but rather its inherent vulnerability. In Fang-Krum, malicious updates are deliberately far from the global model, allowing FLTrust to identify discrepancies easily. In contrast, the malicious updates in Fang-Med are subtler and only slightly deviate from the global model or median, making it harder for FLTrust to distinguish them from benign updates. Meanwhile, FedCC consistently achieves the highest accuracy (89.66%, 70.52%, and 17.83% for each dataset), demonstrating its capability to mitigate attacks involving more nuanced changes to model updates without extreme outliers.
6.2.2 Targeted Backdoor Attacks
Table 5: Test Accuracy under targeted attacks in IID setting.
Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
Target IID fM 88.27 86.63 87.03 89.41 89.45 89.59 75.29 90.01
C10 64.68 57.69 71.19 69.85 68.76 68.61 11.09 71.64
C100 13.83 5.72 17.57 13.24 15.11 17.08 1.03 18.61
DBA C10 10.00 35.58 51.40 10.00 16.16 56.69 10.00 58.04
Figure 4: Confidence of Backdoor Task for targeted attacks in IID settings.
Table 5 and Figure 4 present the test accuracy and backdoor confidence under targeted backdoor attacks in IID settings. Similar to the untargeted attack scenario, the general test accuracy in IID settings is higher than in non-IID settings.
Reducing backdoor confidence is as critical as maintaining main task accuracy, as attackers aim to stealthily embed backdoor tasks. A key observation is FedCC’s remarkable ability to reduce backdoor confidence to near zero while preserving high main task accuracy. In contrast, the fluctuating backdoor confidence observed in Multi-Krum and Flare is largely due to the distributed trigger, which confuses their defense mechanisms and hampers their ability to filter out outliers effectively. Other methods, such as Multi-Krum, Flare, and FLTrust, show significant variability in backdoor confidence, further highlighting FedCC’s superiority.
Notably, FedCC achieved the highest main task accuracy across both targeted backdoor attacks and distributed backdoor attacks (DBA), consistently neutralizing the backdoor task. These results underscore FedCC’s ability to mitigate targeted backdoor attacks through precise and efficient integration of prior knowledge to preserve main task performance while zeroing out the attack confidence.
6.3 Performance in Varying FL Settings
@@
We additionally measured the effectiveness of FedCC in various non-IID settings, including different numbers of malicious clients, fractions of participation, and numbers of local epochs. Since Coomed was as effective as ours, and FLARE also used PLRs, we compared three methods to FedCC, including FedAvg.
6.3.1 Impact of the Number of Malicious Clients
@@
Figure 5: Test Accuracy with different numbers of malicious clients
@@
We assessed the impact of different numbers of malicious clients under untargeted attacks, specifically Untargeted-Krum and Untargeted-Med. In this set of experiments, we kept the total number of clients and participation fraction fixed at 10 and 1, respectively. To adhere to the assumption that the number of attackers is less than half of the participating clients, we investigated the test accuracy considering scenarios involving up to four malicious clients.
@@
Figure 5 illustrates the accuracy in the given experimental settings, and it is evident that FedCC outperforms the other defense methods. We observe a general trend where the accuracy of defense methods decreases as the number of attackers increases. Notably, Coomed experiences the most significant drop in accuracy when transitioning from one malicious client to four malicious clients. This can be attributed to the fact that as the participation of malicious clients increases, there is a higher likelihood of the median values being influenced by their malicious contributions.
@@
6.3.2 Impact of Fraction
Table 6: Test Accuracy under untargeted attacks with different fractions of clients being selected.
Untargeted-Krum Untargeted-Med
@@ 6.3.2 Impact of Fraction
We examine the impact of the participating fraction of clients where the number of total clients is fixed at 100. The fractions are 0.1, 0.3, and 0.5, such that the numbers of clients being aggregated are 10, 30, and 50, respectively. The proportion of malicious clients is fixed at 0.3, such that the malicious clients are 3, 9, and 15, respectively. Table 6 summarizes the accuracy. We can observe that the accuracy tends to increase as the participation fraction is greater under untargeted-Krum attacks. It implies that the more client participation in a non-IID setting, the more accurate the global model is if the attacks are mitigated. Under untargeted-Med attacks, the accuracy is the highest with FedCC and second highest with Coomed. Throughout the experiment, FLARE prevented the global model from convergence, and the test accuracy fluctuated. The fluctuation is presumably due to the weight scaling while randomly selecting clients. Different from previous experiments, the server chooses a fraction of the client every round, meaning not-yet-trained models can be selected. In this environment, weight scaling leveraged by FLARE constantly improperly scales the local weights, causing divergence and fluctuation. These experimental results indicate that selectively choosing clients based on CKA similarity is more reliable than taking the median value of each coordinate in a non-IID setting. We also observed that as the participation fraction grows, the accuracy grows. This is due to the rising number of aggregated local models; the more models are aggregated, the more general global model is generated.
7 Conclusion and Future Work
FL has emerged in response to the rising concerns about privacy breaches while employing AI techniques. FL, however, is vulnerable to model poisoning attacks due to a server’s blindness to local datasets, making it difficult to assume data distribution and model parameter integrity. In line with this, many robust aggregation algorithms and defenses have been proposed; still, existing approaches are orthogonal to various attacks and lack consideration of non-IIDness. Throughout the exhaustive experiments, FedCC mitigates both untargeted and targeted (or backdoor) attacks while demonstrating its effectiveness in non-IID data environments. We leave theoretical guarantees of FedCC and experiments when clients have different neural network architectures or tasks as future work.
@@ 7 Conclusion and Future Work
Appendix 0.A Dataset and Model Architecture
Appendix 0.B FedCC: Algorithm
Algorithm 1 summarizes FedCC’s workflow.
Algorithm 1 FedCC
Input: global_w, n local_w ($w_{1}$, …, $w_{n}$)
      Output: agg_w, larger_cluster_members
$\triangleright$ Get PLRs of local models
1:for $i<n$ do
2:     local_plrs[$i$] $\leftarrow$ local_w[$i$][second_last_layer]
3:end $\triangleright$ Extract global PLR
4:glob_plr $\leftarrow$ global_w[second_last_layer] $\triangleright$ Apply kernel CKA to the plrs
5:cka[$i$] $\leftarrow$ kernel_CKA(glob_plr, local_plrs[$i$]) $\triangleright$ Normalize similarities globally
6:across_cka $\leftarrow$ normalize(cka) $\triangleright$ Apply Kmeans clustering algorithm
7:kmeans $\leftarrow$ kmeans(n_clusters=2, cka)
8:labels $\leftarrow$ kmeans.labels_
9:count $\leftarrow$ counter(labels)
10:larger_cluster $\leftarrow$ 1
11:if count[00] $>$ count[$1$then
12:     larger_cluster $\leftarrow$ 0
13:end $\triangleright$ Identify larger cluster members
14:larger_cluster_members = where(labels == larger_cluster) $\triangleright$ Normalize similarities within larger cluster
15:within_cka $\leftarrow$ normalize(cka[larger_cluster_members])
16:Initialize agg_weights as zero tensor $\triangleright$ Differently weigh weights per layer
17:for layer in model_layers do
18:     if layer is up to second-to-last layer then
19:         agg_weights[layer] $\leftarrow$ weighted average using across_cka
20:     else if layer is second-to-last layer then
21:         agg_weights[layer] $\leftarrow$ weighted average using within_cka
22:     else
23:         agg_weights[layer] $\leftarrow$ weighted average using across_cka
24:     end if
25:end for
26:return agg_weights, larger_cluster_members